DBBC Service - Replaced PCI7200 card, swapped CoMos B & C, reseated stack, connected longer VSI cable to mk5-2ke. No further problems with errorenous readback of power levels, but seem to be suffering a new poblem with clock breaks in the mark5 units. Will revisit earthing of the mark5 units to attempt to correct this problem.
mk5-2ke - Upgraded firmware and software to v3.0.2. Connected to VSI output of DBBC. All working, but still needs an additional counter to fully support its use in recording.
Monica - Rxmon & IFBox code updated to read the temperature & Humidity sensors in control room, maser room, pedestal and hub and the maser battery voltage. Feed temp & RH sensor was damaged in installation but monitor points have been created for when it is available.
UPSMon - A new PC has been installed which monitors the UPS state via NUT, and can send SMSs via an attached mobile phone. This should give warning of any significant power outages.
Phase CAL - The rapid phase wrapping noted by Brian Corey (Dec 2013) appears to be caused by the 5 MHz filter put in place last year. Removing it reduced the phase rate from 360 degree drifts in ~30 min at S-band (~5 min at X-band), to ~30 degrees over 3 hours at S-band (~120-150 degrees over 3.5 hours at X-band). No delay-like differential variations noted in the pcal phases after removing the filter.
Tsys - SLCP LNA appears to be worse than nominal. Adjusted flange connections of S-band feed while monitoring Tsys under load and have got an apparent improvement in SRCP Tsys (~100-110 K instead of ~130 K). New CAL factors have been measured for the cont CAL per band.
pcfske - Discovered that disk 2 of the RAID had failed and that the spare was not in use. Removed disk 2 from the RAID array and set the RAID rebuilding with sda & sdc.
webcamke - Removed webcamke from housing, left housing on pole to avoid a fight with the wasps. A new housing and camera installed further down the pole, replacing webcamke.
/home/www/auscope/opswiki/data/pages/blog/katherine_maintenance_trip_april_12-17.txt · Last modified: 2014/04/24 00:25 by Jamie McCallum